OpenStudy (anonymous):

Help please how do you solve A numbers one digit is 2 more than twice its tens digit. The sum of its digits are 5. What the number?

OpenStudy (unklerhaukus):

let the number be nm, where n is the tens digit, and m is the ones digit

OpenStudy (unklerhaukus):

\[m>2n\]\[m+n=5\]

OpenStudy (anonymous):

Ok thanks

OpenStudy (unklerhaukus):

solve the second equation for m and plug into the first equation,

OpenStudy (anonymous):

How do you do that

OpenStudy (unklerhaukus):

find and inequality for n, remember that n,m are digits, so they must be natural numbers , deduce n, , then find m with the second equation, i have found two solutions

OpenStudy (unklerhaukus):

to solve the second equation for m m+n=5 take n from both sides m=5-n

OpenStudy (anonymous):

I'm sorry I don't really get it

OpenStudy (unklerhaukus):

how can i help ?